### v1

@@ Abstract
This paper demonstrates the feasibility of democratizing AI-driven global weather forecasting models among university research groups by leveraging Graphics Processing Units (GPUs) and freely available AI models, such as NVIDIA’s FourCastNetv2. FourCastNetv2 is an NVIDIA’s advanced neural network for weather prediction and is trained on a 73-channel subset of the European Centre for Medium-Range Weather Forecasts (ECMWF) Reanalysis v5 (ERA5) dataset at single levels and different pressure levels. Although the training specifications for FourCastNetv2 are not released to the public, the training documentation of the model’s first generation, FourCastNet, is available to all users. The training had 64 A100 GPUs and took 16 hours to complete. Although NVIDIA’s models offer significant reductions in both time and cost compared to traditional Numerical Weather Prediction (NWP), reproducing published forecasting results presents ongoing challenges for resource-constrained university research groups with limited GPU availability. We demonstrate both (i) leveraging FourCastNetv2 to create predictions through the designated application programming interface (API) and (ii) utilizing NVIDIA hardware to train the original FourCastNet model. Further, this paper demonstrates the capabilities and limitations of NVIDIA A100’s for resource-limited research groups in universities. We also explore data management, training efficiency, and model validation, highlighting the advantages and challenges of using limited high-performance computing resources. Consequently, this paper and its corresponding GitHub materials may serve as an initial guide for other university research groups and courses related to machine learning, climate science, and data science to develop research and education programs on AI weather forecasting, and hence help democratize the AI NWP in the digital economy.
Iman Khadir1, Shane Stevenson1, Henry Li2, Kyle Krick2, Abram Burrows1, David Hall3, Stan Posey3,
Samuel S.P. Shen1,***Corresponding Author: Samuel S.P. Shen at sshen@sdsu.edu
Department of Mathematics and Statistics, San Diego State University, San Diego, CA 92182, USA
Research & Cyberinfrastructure Division, San Diego State University, San Diego, CA 92182, USA
NVIDIA Corporation, Santa Clara, CA 95051, USA
1 Introduction
@@
The rapid and notable advancement of AI in weather forecasting has dramatically changed the situation. Modern GPUs, such as NVIDIA’s A100 and H200 are affordable by universities and can act as supercomputers on which AI algorithms can run by university research groups. The trend of this remarkable transition of high-resolution numerical weather forecasts from major computing centers to university research labs, or even to university classrooms, has provided unprecedented opportunities for higher education. In particular, both graduate and undergraduate students can practice operational numerical forecasting in their research or coursework. However, a gap exists between the accessibility and usability of advanced AI weather forecasting tools like NVIDIA’s FourCastNet (Pathak et al. 2022) and DeepMind’s GenCast (Lam et al. 2023) for university researchers with limited resources and/or prior AI expertise. The purpose of this paper is to fill this gap by providing users’ guidance for university research labs to run or even to train AI weather forecasting models. We are writing this paper using NVIDIA’s FourCastNetv2 as an example.
@@
NVIDIA is a leader in HPC system development and software technology to support cutting-edge AI weather forecasting. The initial release of FourCastNet (Pathak et al. 2022) was the first major model of this kind to be published. Pathak et al. (2022) was the first major model of this kind to be published. This is a data-driven forecast that uses an Adaptive Neural Fourier Operator (Guibas et al. 2022). Many other AI weather forecasting models have appeared since then, such as Google’s GraphCast (Lam et al. 2023), GenCast (Price et al. 2024), Microsoft’s Aurora (Bodnar et al. 2024), and Huawei’s Pangu-Weather (Bi et al. 2023). Each model has its own specific features and merits. In 2023, NVIDIA released FourCastNetv2, which has a distinct feature of using a Spherical Fourier Neural Operator (Bonev et al. 2023). This updated Version-2 of FourCastNet (FourCastNetv2) improved model stability and accuracy.
@@
Our paper is written based upon FourCastNetv2 for the purpose of tutoring university research groups to make more AI weather forecasts. It is not our intention to evaluate the contributions of each of these AI forecasting models. Instead, we are providing (i) a step-by-step guide to other groups and universities to train students to make AI weather forecasts, and (ii) a suite of Python code to visualize the forecasting products. Our products include visualizations of different atmospheric variables over the globe, comparisons of model forecasts with European Center for Medium-Range Weather Forecasts (ECMWF) Reanalysis 5 (ERA5) data, and the prediction of hurricane tracks. Our work will allow students and instructors to have opportunities to use AI weather forecasting models in university classrooms. This will be an extremely valuable learning experience for students. Specifically, our paper includes examples of 10-day forecasting zonal velocity field at 10 meters vertical height (U10) and a 3-day forecasting for hurricane track for Hurricane Florence from 13 September 2018 to 16 September 2018.
@@ 2.1 Hardware Setup
Running FourCastNet on powerful GPUs requires a nontrivial setup of relevant hardware. This is a major challenge for leveraging large AI models. This challenge should not be underestimated by aspiring university research groups. The AI weather forecasting groups led by a Principal Investigator should build a work connection with the university’s hardware IT team. Figure 1 shows a comprehensive possible interaction among the students and AI research groups with the IT team and AI specialists, such as NVIDIA’s FourCastNet developers.
@@
Figure 2 shows the configuration of the hardware for our research. Through the NSF-funded Technology Infrastructure for Data Exploration (TIDE) project (NSF #2346701), part of the National Research Platform (NRP) Nautilus hyper-cluster, the San Diego State University (SDSU) Climate Informatics Lab has access to up to 4 A100 and 68 L40 NVIDIA GPUs at a time. SDSU’s research IT team has leveraged Kubernetes to optimize the utilization of hardware resources, allowing researchers to easily instantiate and access isolated computing environments within the cluster. Anyone with access can instantiate and gain access to an isolated pod within the cluster. The pods are configured with the packages necessary to interface the hardware with FourCastNet and FourCastNetv2. The usage of isolated pods ensures that each researcher operates in a self-contained environment, preventing interference and resource contention from other users. Moreover, each pod can be configured to use different or multiple GPUs, allowing researchers to dynamically scale projects up or down. Jupyterlab was also installed and configured to run inside each pod so that researchers can quickly manage, edit, and run their code without having to learn Linux terminal commands. The creation of an isolated pod and the configuration of a Jupyterlab allows for new research teams or students to quickly deploy a working environment without the tedious task of chasing down dependencies and, therefore, accelerating the speed of research.
@@
TIDE provides San Diego State University researchers with free access to research-class computational resources focused on AI and machine learning (ML). TIDE offers 17 GPU nodes, each equipped with 4 NVIDIA L40 48 GB VRAM GPUs for a total of 68 L40 GPUs. In addition to the GPU nodes, TIDE has one GPU-Advanced node equipped with 4 NVIDIA A100 80 GB VRAM GPUs. For workloads that do not require access to GPUs, TIDE includes 6 CPU nodes with 64 CPUs, each totaling 384 CPUs. The CPU nodes are also equipped with higher amounts of RAM to support CPU-only workloads. TIDE has three storage nodes that are configured as a Ceph cluster with triple replication and offers a combined total of 240 TB usable storage. TIDE leverages a 100 Gbps HPR connection via Science DMZ to facilitate research network connectivity with other institutions. TIDE also has a connection to the public internet for accessing and publishing public datasets. TIDE is integrated into the National Research Platform (NRP) Nautilus distributed Kubernetes cluster. Kubernetes provides management of storage, RAM, CPU, and GPU requests. Code is packaged into software containers, which are then run in Kubernetes via both interactive and batch jobs.
@@ 2.2 Training Data
FourCastNet and FourCastNetv2 both utilize the ERA5 dataset for training, but they differ in the number of variables selected for inclusion. Specifically, FourCastNet uses a subset of 20 variables, while FourCastNetv2 leverages a broader subset of 73 variables. Both models use a 6-hour subsample of the hourly ERA5 dataset.
@@ 2.3 Storage of the Training Data
Approximately 6 terabytes in size, the ERA5 Reanalysis dataset was downloaded from the Copernicus Climate Store by the San Diego State University IT team using Globus. The data is stored in a shared CephFS volume inside the TIDE cluster, ensuring accessibility for all project researchers. Each researcher has their own pod with different computation resources; each pod has access to this same shared volume and data. Within such a shared volume are also NVIDIA’s model parameters as well as key statistical measures of the data, which have been precomputed from ERA5.
@@ 2.4 Initialization of the forecast
The initialization of forecasts using FourCastNetv2 and FourCastNet begins by accessing a single time step of the required variables in the ERA5 dataset. This can be achieved in many ways, including the Climate Data Store (CDS) API, ECMWF Web Interface, Amazon Web Services (AWS), Google Earth Engine, or simply accessing local files given that the dataset is downloaded locally. For this project, we chose to use the CDS API and our local copy of the ERA5 dataset.
To access the CDS API, users must first create an account, install the CDS API, and place the provided authentication key in the designated location on their local device’s file system. Once access to the API is successfully established, users can query it to download the necessary variables. Detailed instructions and code to facilitate this process are available in our user’s manual on Github (Khadir and Stevenson 2025). After the required data has been downloaded, the next step is to arrange the required variables into data arrays with the specific dimensions for the AI forecasting models to process. For FourCastNetv2, the input data array must have the dimension , while for FourCastNet, the required data array dimension is . The first dimension corresponds to the set of variables utilized by each model during inference, and the latter two dimensions represent the 0.25° spatial resolution of the ERA5 dataset.
Once the dataset has been gathered and organized into the correct dimensions, the data must be normalized before it is given to the model. This step involves normalizing the input data to ensure that all variables have a mean of and a standard deviation of . This normalization is performed using the formula:
@@
After each of the variables has been normalized, FourCastNetv2 requires a different process than FourCastNet to produce the final forecasting products. For FourCastNet, the model needs to be instantiated using the model’s source code. Once initialized, users should pass the normalized dataset, the index of the variable they wish to forecast, the desired prediction length, and the model itself into NVIDIA’s autoregressive influencer pipeline. This then generates the forecast by iteratively using model output to initialize new predictions. The code required to accomplish this step is documented in detail in our user’s manual (Khadir and Stevenson 2025).
For FourCastNetv2, users can access the model through the ECMWF Experimental AI-Models API, rather than working directly with the model’s source code. This API simplifies the process by automatically handling the normalization of input data, eliminating the need for users to perform this step manually. Additionally, users can integrate the ECMWF API with their CDS credentials, allowing the AI-Models API to handle the data download process as well. A comprehensive guide to setting up and using the API is provided in our user’s manual.
@@ 3 Results
To show practical AI forecasting scenarios, we chose to present two cases, one is a three-day hurricane track prediction, and another is on 24-hour and 150-hour predictions of two meteorological variables: zonal wind at 10 m height and surface atmospheric pressure.
@@ 3.1 Predictions of Hurricane Florence
The predictants are mean sea level pressures (SLP) for a 102-hour period starting from September 13th, 2018, at 12:00 AM, with predictions made in 6-hour intervals. After generating the predictions, we identify the eye of the hurricane by locating the minimum mean SLP at each 6-hour interval.
Figure 3 presents two different predictions of hurricane dynamics from two different models: FourCastNet (a) and FourCastNetv2 (b). The initial conditions are shown by the background color map for mean SLP and by the black arrows for the wind field at 12:00 AM, September 13th, 2018. The predictions are the locations of the hurricane eyes, defined as the location of minimum sea level pressure for a given time and represented by the red dots for predictions compared with the pink dots for the ground truth.

### v3

@@ Abstract
This paper demonstrates the feasibility of democratizing AI-driven global weather forecasting models among university research groups by leveraging Graphics Processing Units (GPUs) and freely available AI models, such as NVIDIA’s FourCastNetv2. FourCastNetv2 is an NVIDIA’s advanced neural network for weather prediction and is trained on a 73-channel subset of the European Centre for Medium-Range Weather Forecasts (ECMWF) Reanalysis v5 (ERA5) dataset at single levels and different pressure levels. Although the training specifications for FourCastNetv2 are not released to the public, the training documentation of the model’s first generation, FourCastNet, is available to all users. The training had 64 A100 GPUs and took 16 hours to complete. Although NVIDIA’s models offer significant reductions in both time and cost compared to traditional Numerical Weather Prediction (NWP), reproducing published forecasting results presents ongoing challenges for resource-constrained university research groups with limited GPU availability. We demonstrate both (i) leveraging FourCastNetv2 to create predictions through the designated application programming interface (API), and (ii) utilizing NVIDIA hardware to train the original FourCastNet model. Further, this paper demonstrates the capabilities and limitations of NVIDIA A100’s for resource-limited research groups in universities. We also explore data management, training efficiency, and model validation, highlighting the advantages and challenges of using limited high-performance computing resources. Consequently, this paper and its corresponding GitHub materials may serve as an initial guide for other university research groups and courses related to machine learning, climate science, and data science to develop research and education programs on AI weather forecasting, and hence help democratize AI NWP in the digital economy.
@@
The rapid and notable advancement of AI in weather forecasting has dramatically changed the situation. Modern GPUs, such as NVIDIA’s A100 and H200 are affordable by universities and can act as supercomputers on which AI algorithms can run by university research groups. The trend of this remarkable transition of high-resolution numerical weather forecasts from major computing centers to university research labs, or even to university classrooms, has provided unprecedented opportunities for higher education. In particular, both graduate and undergraduate students can practice operational numerical forecasting in their research or coursework. However, a gap exists between the accessibility and usability of advanced AI weather forecasting tools like NVIDIA’s FourCastNet (Pathak et al., 2022) and DeepMind’s GenCast (Lam et al., 2023) for university researchers with limited resources and/or prior AI expertise. The purpose of this paper is to fill this gap by providing users’ guidance for university research labs to run or even to train AI weather forecasting models. We are writing this paper using NVIDIA’s FourCastNetv2 as an example.
@@
NVIDIA is a leader in HPC system development and software technologies that support cutting-edge AI weather forecasting. The initial release of FourCastNet (Pathak et al., 2022) was the first major model of this kind to be published. This is a data-driven forecast that uses an Adaptive Neural Fourier Operator (Guibas et al., 2022). Many other AI weather forecasting models have appeared since then, such as Google’s GraphCast (Lam et al., 2023), GenCast (Price et al., 2024), Microsoft’s Aurora (Bodnar et al., 2024), and Huawei’s Pangu-Weather (Bi et al., 2023). Each model has its own specific features and merits. In 2023, NVIDIA released FourCastNetv2, which has a distinct feature of using a Spherical Fourier Neural Operator (Bonev et al., 2023). This updated Version-2 of FourCastNet (FourCastNetv2) improved model stability and accuracy. FourCastNet is part of NVIDIA’s efforts to create a digital twin of the Earth through an accessible platform called Earth-2. Earth-2 weather analytics brings together the latest HPC climate technologies, which will allow for scientists, governments, and enthusiasts to better prepare for weather and climate related disasters (Edwards, 2025). An example of the visualization tools Earth-2 Weather Analytics provides is a graphical user interface, where users can generate predictions using FourCastNetv2 and view them on an interactive, 3D globe that can be rotated, zoomed, and explored from any angle (NVIDIA, 2025).
@@
Our paper is written based upon FourCastNetv2 for the purpose of tutoring university research groups to make more AI weather forecasts. It is not our intention to evaluate the contributions of each of these AI forecasting models. Instead, we are providing (i) a step-by-step guide to other groups and universities to train students to make AI weather forecasts, and (ii) a suite of Python code to visualize the forecasting products. Our products include global visualizations of different atmospheric variables, comparisons of model forecasts with European Center for Medium-Range Weather Forecasts (ECMWF) Reanalysis 5 (ERA5) hourly data (Hersbach et al., 2023), and prediction of hurricane tracks. Our work will allow students and instructors to have opportunities to use AI weather forecasting models in university classrooms. This will be an extremely valuable learning experience for students. Specifically, our paper includes examples of 10-day forecasting zonal velocity field at 10 meters vertical height (U10) and a 3-day track for Hurricane Florence from 13 September 2018 to 16 September 2018.
@@ 1.1 Hardware Setup
Running FourCastNet on powerful GPUs requires nontrivial setup of relevant hardware. This is a major challenge for leveraging large AI models. This challenge should not be underestimated by aspiring university research groups. The AI weather forecasting groups led by a Principal Investigator should build a work connection with the university’s hardware IT team. Figure 1 shows a comprehensive possible interaction among the students and AI research groups with the IT team and AI specialists, such as NVIDIA’s FourCastNet developers.
@@
Figure 2 shows the configuration of the hardware for our research. Through the NSF-funded Technology Infrastructure for Data Exploration (TIDE) project, a part of the National Research Platform (NRP) Nautilus hyper-cluster, the San Diego State University (SDSU) Climate Informatics Lab has access up to 4 A100 and 68 L40 NVIDIA GPUs at a time. SDSU’s research IT team has leveraged Kubernetes to optimize the utilization of hardware resources, allowing researchers to easily instantiate and access isolated computing environments within the cluster. Kubernetes is an open-source system for automating the deployment and scaling of applications. Anyone with access can instantiate and gain access to an isolated pod within the cluster. The pods are configured with the packages necessary to interface the hardware with FourCastNet and FourCastNetv2. The usage of isolated pods ensures that each researcher operates in a self-contained environment, preventing interference and resource contention from other users. Moreover, each pod can be configured to use different or multiple GPUs allowing researchers to dynamically scale projects up or down. Jupyterlab was also installed and configured to run inside of each pod so that researchers can quickly manage, edit, and run their code without having to learn Linux terminal commands. The creation of an isolated pod and the configuration of a Jupyterlab allows for new research teams or students to quickly deploy a working environment without the tedious task of chasing down dependencies and, therefore, accelerating the speed of research.
@@
TIDE provides San Diego State University researchers with free access to research-class computational resources focused on AI and machine learning (ML). TIDE offers 17 GPU nodes each equipped with 4 NVIDIA L40 48 GB VRAM GPUs for a total of 68 L40 GPUs. In addition to the GPU nodes, TIDE has one GPU-Advanced node equipped with 4 NVIDIA A100 80 GB VRAM GPUs. For workloads that do not require access to GPUs, TIDE includes 6 CPU nodes with 64 CPUs each totaling 384 CPUs. The CPU nodes are also equipped with higher amounts of RAM to support CPU-only workloads. TIDE has 3 storage nodes that are configured as a Ceph cluster with triple replication and offers a combined total of 240 TB usable storage. TIDE leverages a 100Gbps HPR connection via Science DMZ to facilitate research network connectivity with other institutions. TIDE also has a connection to the public internet for accessing and publishing public datasets. TIDE is integrated into the National Research Platform (NRP) Nautilus distributed Kubernetes cluster. Kubernetes provides management of storage, RAM, CPU, and GPU requests. Code is packaged into software containers which are then run in Kubernetes via both interactive and batch jobs.
@@ 1.2 Training Data
FourCastNet and FourCastNetv2 both utilize the ERA5 dataset for training, but they differ in the number of variables selected for inclusion. Specifically, FourCastNet uses a subset of 20 variables, while FourCastNetv2 leverages a broader subset of 73 variables. Both models use a 6 hour subsample of the hourly ERA5 dataset.
@@ 1.3 Storage of the Training Data
Approximately 6 terabytes in size, the ERA5 Reanalysis dataset was downloaded from the Copernicus Climate Store by the San Diego State University IT team using Globus. The data is stored in a shared CephFS volume—a distributed, POSIX-compliant file system provided by the Ceph storage platform—inside the TIDE cluster, ensuring accessibility for all project researchers. Each researcher has their own pod with different computation resources, each pod has access to this same shared volume and data. Within such a shared volume are also NVIDIA’s model parameters as well as key statistical measures of the data which have been precomputed from ERA5.
@@ 1.4 Initialization of the forecast
The initialization of forecasts using FourCastNetv2 and FourCastNet begins by accessing a single time step of the required variables in the ERA5 dataset. This can be achieved in many ways, including the Climate Data Store (CDS) API, ECMWF Web Interface, Amazon Web Services (AWS), Google Earth Engine, or simply accessing local files given the dataset is downloaded locally. For this project, we chose to use the CDS API and our local copy of the ERA5 dataset.
To access the CDS API, users must first create an account, install the CDS API, and place the provided authentication key in the designated location on their local device’s file system. Once access to the API is successfully established, users can query it to download the necessary variables. Detailed instructions and code to facilitate this process are available in our user’s manual on Github (Khadir and Stevenson, 2025). After the required data has been downloaded, the next step is to arrange the required variables into data arrays with the specific dimensions for the AI forecasting models to process. For FourCastNetv2, the input data array must have the dimension , while for FourCastNet, the required data array dimension is . The first dimension corresponds to the set of variables utilized by each model during inference, and the latter two dimensions represent the 0.25° spatial resolution of the ERA5 dataset.
Once the dataset has been gathered and organized into the correct dimensions, the data must be normalized before it is given to the model. This step involves normalizing the input data to ensure that all variables have a mean of and a standard deviation of . This normalization is performed using the following formula:
@@
After each of the variables has been normalized, FourCastNetv2 requires a different process than FourCastNet to produce the final forecasting products. For FourCastNet, the model needs to be instantiated using the model’s source code. Once initialized, users should pass the normalized dataset, the index of the variable they wish to forecast, the desired prediction length, and the model itself into NVIDIA’s autoregressive influencer pipeline. This then generates the forecast by iteratively using model output to initialize new predictions. The code required to accomplish this step is documented in detail in our user’s manual (Khadir and Stevenson, 2025).
For FourCastNetv2, users can access the model through the ECMWF Experimental AI-Models API (Raoult et al., 2022), rather than working directly with the model’s source code. This API simplifies the process by automatically handling the normalization of input data, eliminating the need for users to perform this step manually. Additionally, users can integrate the ECMWF API with their CDS credentials, allowing the AI-Models API to handle the data download process as well. A comprehensive guide to setting up and using the API is provided in our user’s manual.
@@ 2 Results
To show practical AI forecasting scenarios, we chose to present two cases, one is a three-day hurricane track prediction, and another is on 24-hour and 150-hour predictions of two meteorological variables: zonal wind at 10 m height, and surface atmospheric pressure.
@@ 2.1 Predictions of Hurricane Florence
2.1.1 Model initialization and validation data
(a) Data for the initial conditions
Forecast initialization with FourCastNet and FourCastNetv2 requires a global one-hour mean dataset comprising all 20 variables for FourCastNet and 73 variables for FourCastNetv2. All input variables are needed when to perform a forecast. By default, the forecast is for every variable. Namely, when running FourCastNetv2, all 73 variables are being predicted. A user may choose to output the forecast result for one or more variables, such as mean sea level pressure and 10-meter zonal wind. Table 1 lists the 73 required variables of FourCastNetv2.
Table1. FourCastNetv2’s 73 variables.
| Variable | Description | Pressure level (hPa) |
| u10 | 10 meter u-wind component | – |
| u100 | 100 meter u-wind component | – |
| v10 | 10 meter v-wind component | – |
| v100 | 100 meter v-wind component | – |
| t2 | 2 meter temperature | – |
| sp | Surface pressure | – |
| msl | Mean sea level pressure | – |
| tcwv | Total column vertically-integrated water vapor | – |
| z— | Geopotential (at pressure level —) | 50, 100, 150, 200, 250, 300, 400, 500, 600, 700, 850, 925, 1000 |
| t— | Temperature (at pressure level —) | |
| u— | u component of the wind (at pressure level —) | |
| v— | v component of the wind (at pressure level —) | |
| r— | Relative humidity (at pressure level —) |
Table 1’s first eight variables, such as 10-meter zonal u–wind component and mean sea level pressure (SLP), in the upper half of the table are variables on a single layer of the atmosphere. The five parameters, such as temperature and relative humidity, are for 13 layers of the atmosphere, divided according to the atmospheric pressure levels, 50, 100, , 1,000 hPa. These five parameters at the 13 layers are counted as 65 variables (). Thus, altogether we have 73, which is 8 plus 65, variables for FourCastNetv2.
(b) Validation design
We chose to use Hurricane Florence as an example in this paper. Hurricane Florence occurred in 2018, which is the year designated for validation and therefore excluded from the training sets of both FourCastNet and FourCastNetv2. This ensures that the models have not previously “seen” the event during training, allowing for an objective evaluation of their forecasting accuracy. This is the same reason why 2018’s Hurricane Michael was originally selected by Pathak et al. (2022) for their forecast validation.
Our validation will be focused on the hurricane track. The predictants are mean sea level pressures (SLP) for a 102-hour period starting from September 13th, 2018, at 12:00 AM, with predictions made in 6-hour intervals. After generating the predictions, we identify the eye of the hurricane by locating the minimum mean SLP at each 6-hour interval.
2.1.2 Prediction Results
Figure 3 presents two different predictions of hurricane dynamics from two different models: Fig. 3 (a) for FourCastNet, and Fig. 3(b) FourCastNetv2. The initial conditions are shown by the background color map for mean SLP and by the black arrows for the wind field at 12:00 AM, September 13th, 2018. The mean SLP field at 12:00 AM, September 13th, 2018 is the initial condition for one of the 73 variables. We use the gridded lat-lon hourly ECMWF ERA5 data (Hersbach et al., 2023) for all 73 variables at 12:00 AM, September 13th, 2018, as the complete initial condition to make our FourCastNetv2 predictions.
The predictions are the locations of the hurricane eyes, which are defined as the locations of minimum sea level pressure at a given time and represented by the red dots for predictions. The predicted hurricane track, for approximately four days from 12:00 AM, September 13 to 18:00, September 16, is compared with the ground truth represented by the pink dots. See Fig. 3.